# First Astrometric Limits on Binary Planets and Exomoons orbiting Pictoris b

### v1

@@ 1 Introduction
As thousands of exoplanets have been discovered in the past 30 years, interest in whether moons orbit these worlds has grown. Despite the prevalence of moons in the Solar System (over 400 known111https://ssd.jpl.nasa.gov/sats/discovery.html) and the detection of a few candidates outside the Solar System (e.g. Teachey & Kipping, 2018; Lazzoni et al., 2020; Kipping et al., 2022), no extrasolar moons, or exomoons, have been confirmed. Nevertheless, given the well-established moon formation pathways benchmarked in the Solar System, theoretical studies suggest that exomoons should exist. One dominant mechanism in the Solar System is co-formation with the planets in circumplanetary disks, where moons form alongside their host planets in a gas-rich disk, similar to the Galilean moons of Jupiter (Canup & Ward, 2006). Another formation pathway is capture events, where a passing object becomes gravitationally bound to a planet. This process has been proposed as a key mechanism for the formation of moons on eccentric, inclined, and often retrograde orbits in our Solar System (Hamers & Portegies Zwart, 2018), including Triton (Agnor & Hamilton, 2006). Furthermore, debris generated from massive impacts, such as the hypothesized origin of the Earth-Moon system, could also lead to exomoon formation (Barr & Bruck Syal, 2017).
@@
The hunt for exomoons is especially important for understanding the formation and evolution of planetary systems. Just as the Moon plays a key role in Earth’s orbital stability and tides, exomoons are expected to gravitationally influence their host planets, potentially affecting their rotational dynamics and long-term climate stability (Laskar et al., 1993; Poon et al., 2024). The successful detection of an exomoon also has the potential to transform our understanding of migration theories and lunar habitability (Barnes & O’Brien, 2002; Sucerquia et al., 2019, 2020; Martínez-Rodríguez et al., 2019).
@@
Despite the importance of discovering exomoons, their detection remains extremely challenging. The signals from moons are often too subtle and faint to detect directly as a result of the limited resolution and sensitivity of current telescopes. Indirect methods, such as analyzing light curve variations during planetary transits, offer alternative ways to detect exomoons (Kipping et al., 2015; Teachey & Kipping, 2018; Kipping, 2021). Researchers search for signals such as transit timing variations (TTVs, Kipping, 2009, 2020a, 2020b), transit duration variations (TDVs, Kipping et al., 2009), and transit radius variations (TRVs, Rodenbeck et al., 2020). From these methods, a handful of exomoon candidates have been proposed, including Kepler-1625b-i, a potential Neptune-sized moon orbiting a super-Jovian planet (Teachey & Kipping, 2018), and Kepler-1708b-i, a potential 2.6 Earth-radii moon orbiting a Jupiter-sized planet at a distance of 12 planetary radii (Kipping et al., 2022). However, none of these candidates have been confirmed, and the detections are controversial (e.g. Kreidberg et al., 2019; Heller & Hippke, 2024). We note that similar methods can also be used to detect binary planets (Chakraborty & Kipping, 2022), but none have been identified.
@@
With no confirmed exomoons, research has instead focused on placing increasingly strong limits on their possible presence. For instance, a considerable amount of work has gone into placing upper bounds on the masses and radii of exomoons based on transit observations (e.g. Kipping et al., 2013, 2014, 2015; Kipping, 2020a, b). More recently, other detection methods have been used to place constraints, including radial velocity (RV) observations of directly imaged planets (Vanderburg et al., 2018; Vanderburg & Rodriguez, 2021; Ruffio et al., 2023; Horstman et al., 2024). These findings demonstrate that, although the confident detection of exomoons is still elusive, current methods can place constraints on their masses and orbits.
@@
Several studies have suggested that astrometry can be used to identify and constrain lunar companions within a system (Cabrera & Schneider, 2007; Agol et al., 2015; van Woerkom & Kleisioti, 2024). This is a particularly exciting method given the recent advancements in measuring precise astrometry of exoplanets using long-baseline interferometry. Recently, Lacour et al. (2021) showcased the high-precision astrometric capabilities of the GRAVITY instrument (GRAVITY Collaboration et al., 2017) on the Very Large Telescope Interferometer (VLTI), which can make astrometric measurements with precision of tens of microarcseconds. Building on these capabilities, Winterhalder et al. (2025) demonstrated through simulations and VLTI/GRAVITY observations that moons as small as 0.14 could already be inferred around the nearby (d26.9 pc) AF Lep b system, with future facilities expected to reach Earth-like moons in habitable zones. Complementary work by Wagner et al. (2025) presented a simulation framework for detecting exomoons around imaged giant planets, showing that Earth-mass moons in the habitable zone of Centauri A could be detected with future space- or ground-based facilities. Most recently, Kral et al. (2025) placed constraints on moons orbiting the brown dwarf HD 206893 B with GRAVITY astrometry and detected a tentative candidate. Together, these studies highlight how astrometry is emerging as a promising pathway toward the first confirmed detections of exomoons.
@@
One of the most promising targets for astrometric studies is the Beta () Pictoris system, a young and nearby planetary system that has been extensively studied since the discovery of its circumstellar debris disk in the 1980s (Smith & Terrile, 1984). This system hosts two directly imaged giant planets, Pic b and c, making it ideal for testing astrometric detection techniques. Pic b, located at a semi-major axis of AU, was discovered through direct imaging, while Pic c, at AU, was discovered through RV measurements (Lagrange et al., 2010, 2020). Lagrange et al. (2020) combined astrometry and RV data to constrain the orbits and masses of both planets, finding that Pic b has a mass of and Pic c has a well-constrained mass of .
@@
In this work, we use archival data from Lagrange et al. (2020) and Lacour et al. (2020a) to place constraints on the presence of exomoons and binary planets in the Pictoris system using high-precision astrometric measurements from the GRAVITY instrument. While we find no evidence supporting the existence of an exomoon, our analysis establishes strong constraints on the mass and orbital period of any potential companions, providing a framework for applying astrometric techniques to future exomoon searches. Our paper is organized as follows: Section 2 outlines the instrumentation and data. Section 3 describes our analysis methods, including the development and application of our model. Section 4 presents our results, focusing on the derived mass and orbital constraints. Section 5 then discusses the implications of our findings, comparing astrometry with complementary techniques and addressing the potential for future studies. We conclude in Section 6.
2 Observations
@@
In this work, we used existing archival direct imaging and interferometric observations of Pictoris b and c. Pictoris has been observed for over a decade by numerous instruments, and the precision of astrometric measurements has increased over time. We focused on data taken with the GRAVITY interferometer (GRAVITY Collaboration et al., 2017) and the SPHERE high-contrast imager (Beuzit et al., 2019), provided in Table 1. Following Lacour et al. (2020a), we excluded earlier, less-precise data from the NACO (NAOS-CONICA) instrument (Lenzen et al., 2003; Rousset et al., 2003). Together, these datasets provide a high-precision foundation for our orbital analysis.
@@ 2.2 GRAVITY
We use GRAVITY K band observations of Pictoris b and c from Lagrange et al. (2020) and Lacour et al. (2020a). The observations collected by Lagrange et al. (2020) took place on six nights between September 22, 2018, and February 9, 2020. Three of the observations were obtained in the medium-resolution (R = 500) mode of the GRAVITY spectrometer, and two were obtained in the high-resolution mode (R = 4000). On January 7th, both modes were used. Following Lacour et al. (2021), we exclude two of the GRAVITY observations of Lagrange et al. (2020) (MJDs 58738.267 and 58855.198) from our analysis due to poor coherence time and limited exposures. We also include in our analysis additional observations collected by Lacour et al. (2020a) as part of the ExoGRAVITY program (Lacour et al., 2020b) on the night of January 6, 2021 for Pictoris c and the nights of January 7 and August 27, 2021, for Pictoris b. The astrometric calibration from both Lagrange et al. (2020) and Lacour et al. (2020a) underwent processing using version 1.5.0 of the ESO GRAVITY pipeline (Lapeyrere et al., 2014).
3 Analysis
@@ 3.1 Brightness Ratio
3.2 Two Planet Model and Likelihood Function
To fit the SPHERE and GRAVITY data, we constructed a model capable of accurately predicting the positions of Pic b and c over time. To create this initial two-planet model, we adapted code from the Orbitize! Python package (Blunt et al., 2020) to simultaneously calculate the right ascension (RA) and declination (DEC) offsets of both planets with respect to the star. The model also accounts for the astrometric perturbation on Pic b caused by its inner companion, Pic c, following the methodology of Blunt et al. (2020).
We constructed a log likelihood function to compare the archival data (see Section 2) with this model. Maximizing the log likelihood () determines the best-fit orbital parameters for Pic b and c, and exploring the parameter space constrains the plausible parameter space for any orbiting exomoon. We found that the uncertainties on the GRAVITY observations appeared to be slightly underestimated, so we incorporated a multiplicative “jitter” term to allow our fit to adjust the GRAVITY uncertainties such that:
@@ 3.2 Two Planet Model and Likelihood Function
3.3 Two Planet & Moon Model
To model the astrometric signature of a hypothetical dark exomoon orbiting Pic b, we extend the two-planet model described in Section 3.2 by adding a Keplerian companion on a bound orbit around Pic b. We used the same log likelihood functions as described in Equations 1 through 4, but include an additional astrometric perturbation term representing the motion of Pic b about the barycenter of the Pic b-moon system. To do this, we used orbitize to generate Keplerian position offsets (RA, DEC) for an exomoon orbiting Pic b, where its gravitational influence shifts the apparent position of Pic b by:
@@ 3.3 Two Planet & Moon Model
| (M⊙) | 1.701 | 1.702 | 1.73 |
| Parallax (mas) | 51.39 | 51.39 | 51.44 |
| Error Scale Factor | 1.31 | 1.35 | – |
@@
| Semi-Major Axis (au) | 10.03 | 10.04 | 9.95 |
| Eccentricity | 0.1094 | 0.1082 | 0.106 |
| Inclination (deg) | 88.991 | 88.994 | 88.99 |
| Argument of Periastron (deg) | 212.5 | 213.1 | 203.2 |
| Longitude of Ascending Node (deg) | 31.773 | 31.770 | 31.80 |
| Epoch of Periastron Passage | 0.774 | 0.775 | 0.730 |
@@
| Semi-Major Axis (au) | 2.688 | 2.681 | 2.61 |
| Eccentricity | 0.235 | 0.235 | 0.32 |
| Inclination (deg) | 88.94 | 88.94 | 88.92 |
| Argument of Periastron (deg) | 61 | 60 | 60.8 |
| Longitude of Ascending Node (deg) | 31.18 | 31.19 | 31.07 |
| Epoch of Periastron Passage | 0.825 | 0.823 | 0.708 |
3.4 Fitting
@@
4 Results
4.1 Constraints on Exomoons Around Pic b
@@ 4.2 Orbital Parameters for Pic b and c
| Orbital Period (days) | 68% (M) | 95% (M) | 99.7% (M) |
| 2.424749 | 1.180788 | 3.009138 | 4.857555 |
| 7.274247 | 0.800393 | 1.896973 | 3.122645 |
| 12.123746 | 0.480739 | 1.014212 | 1.710106 |
| 16.973244 | 0.342458 | 0.675760 | 1.130730 |
| 21.822742 | 0.280340 | 0.543695 | 0.886558 |
5 Discussion
5.1 Constraints on Exomoons and Binary Planets in Context
@@
The moons that we could detect in the Pic system are not like the moons in our own solar system. The Solar System’s most massive moon is Ganymede, which has a mass 50 times smaller than that of Earth, and 1500 times smaller than the lowest mass moon we can rule out at 3 around Pic b. However, the moons and binary planets that we could detect in the Pic system have similar mass ratios to some of the Solar System moons. At long orbital periods, we could detect moons with masses of about 1%, which is smaller than both the Pluto/Charon and Earth/Moon mass ratios. It is worth noting that both the Earth/Moon and Pluto/Charon systems are believed to have formed via collision and capture of lunar material (Denton et al., 2025; Canup & Asphaug, 2001); any moons we could detect might likely have formed similarly in their systems.
5.2 Are we sensitive to moons or binary planets?
@@
Given that we are only sensitive to Neptune-mass companions and larger, it is worth asking: are the objects that we could detect moons, or should they instead be considered binary planets? The classification of a satellite as a moon or a binary planet remains an open question, particularly when considering systems beyond our solar system. A simple way to distinguish moons and binary planets would be to choose a mass ratio between the two orbiting objects above which an orbiting body is a binary planet, and below which it is a moon. Such a strategy has been applied to the IAU definition of a planet (Lecavelier des Etangs & Lissauer, 2022), which states that the mass ratio must be large enough that stable orbits can exist around the L4 and L5 Lagrange points: , or more precisely . If we were to, by extension, apply this definition to distinguish moons from binary planets, our analysis is sensitive to some exomoons around Pic b at orbital periods longer than about 100 days.
@@
However, there is no official IAU definition for a moon, so instead, other criteria have been utilized. According to one commonly cited definition from Stern & Levison (2002), a moon is a satellite with a mass and semi-major axis small enough to keep the center of mass of the moon-planet system within the radius of the primary planet. In contrast, a binary planet is defined as a system in which the companion is sufficiently massive for the center of mass to extend beyond the primary planet. By this definition, we are not sensitive to exomoons around Pic b. We show the Stern & Levison (2002) boundary between moons and binary planets in Figure 4 as a solid green line. This boundary falls slightly below our 1 contour at all orbital periods and scales in a similar manner with orbital period to our sensitivity limits. This similar scaling is expected, since the Stern & Levison (2002) definition of a moon and the amplitude of astrometric signals both depend on the center of mass motion of the planet.
@@
However, both of these definitions have drawbacks. For instance, although Charon formally meets the requirement for a binary planet from both of these two definitions, it is still commonly referred to as a moon of Pluto. Moreover, under the Stern & Levison (2002) definition, the same object at different orbital separations may be either a moon or a binary planet. Given orbital evolution, this implies that an orbiting companion may be both a moon and a binary planet at different times in the system’s history. For instance, while the Earth/Moon system currently meets the Stern & Levison (2002) criterion for a moon, after the Moon’s semimajor axis has increased by about 40% due to tidal evolution, it would by this definition formally become a binary planet. It is also plausible that because planet radii contract over time, a young planet/moon system may evolve into a binary system as the planet radius becomes small enough that it does not encompass the system’s center of mass. Ultimately, a precise distinction between moons and binary planets will likely have to wait until astronomers have a better understanding of the broader population beyond the solar system.
5.3 Prospects for Improved Exomoon Searches in the Future

### v2

@@ I Introduction
As thousands of exoplanets have been discovered in the past 30 years, interest in whether moons orbit these worlds has grown. Despite the prevalence of moons in the Solar System (over 400 known111https://ssd.jpl.nasa.gov/sats/discovery.html) and the detection of a few candidates outside the Solar System (e.g. Teachey and Kipping, 2018; Lazzoni et al., 2020; Kipping et al., 2022), no extrasolar moons, or exomoons, have been confirmed. Nevertheless, given the well-established moon formation pathways benchmarked in the Solar System, theoretical studies suggest that exomoons should exist. One dominant mechanism in the Solar System is co-formation with the planets in circumplanetary disks, where moons form alongside their host planets in a gas-rich disk, similar to the Galilean moons of Jupiter (Canup and Ward, 2006). Another formation pathway is capture events, where a passing object becomes gravitationally bound to a planet. This process has been proposed as a key mechanism for the formation of moons on eccentric, inclined, and often retrograde orbits in our Solar System (Hamers and Portegies Zwart, 2018), including Triton (Agnor and Hamilton, 2006). Furthermore, debris generated from massive impacts, such as the hypothesized origin of the Earth-Moon system, could also lead to exomoon formation (Barr and Bruck Syal, 2017).
@@
The hunt for exomoons is especially important for understanding the formation and evolution of planetary systems. Just as the Moon plays a key role in Earth’s orbital stability and tides, exomoons are expected to gravitationally influence their host planets, potentially affecting their rotational dynamics and long-term climate stability (Laskar et al., 1993; Poon et al., 2024). The successful detection of an exomoon would also provide valuable constraints on lunar habitability and planetary migration theories, as the survival, orbital architecture, and tidal evolution of moons depend sensitively on how and when their host planets migrate (Barnes and O’Brien, 2002; Sucerquia et al., 2019, 2020; Martínez-Rodríguez et al., 2019).
@@
Despite the importance of discovering exomoons, their detection remains extremely challenging. The signals from moons are often too subtle and faint to detect directly as a result of the limited resolution and sensitivity of current telescopes. Indirect methods, such as analyzing light curve variations during planetary transits, offer alternative ways to detect exomoons (Kipping et al., 2015; Teachey and Kipping, 2018; Kipping, 2021). Researchers search for signals such as transit timing variations (TTVs, Kipping, 2009, 2020a, 2020b), transit duration variations (TDVs, Kipping et al., 2009), and transit radius variations (TRVs, Rodenbeck et al., 2020). From these methods, a handful of exomoon candidates have been proposed, including Kepler-1625b-i, a potential Neptune-sized moon orbiting a super-Jovian planet (Teachey and Kipping, 2018), and Kepler-1708b-i, a potential 2.6 Earth-radii moon orbiting a Jupiter-sized planet at a distance of 12 planetary radii (Kipping et al., 2022). However, none of these candidates have been confirmed, and the detections are controversial (e.g. Kreidberg et al., 2019; Heller and Hippke, 2024). We note that similar methods can also be used to detect binary planets (Chakraborty and Kipping, 2022), but none have been identified.
@@
With no confirmed exomoons, research has instead focused on placing increasingly strong limits on their possible presence. For instance, a considerable amount of work has gone into placing upper bounds on the masses and radii of exomoons based on transit observations (e.g. Kipping et al., 2013, 2014, 2015; Kipping, 2020a, b). More recently, other detection methods have been used to place constraints, including radial velocity (RV) observations of directly imaged planets (Vanderburg et al., 2018; Vanderburg and Rodriguez, 2021; Ruffio et al., 2023; Horstman et al., 2024). These findings demonstrate that, although the confident detection of exomoons is still elusive, current methods can place constraints on their masses and orbits.
@@
Several studies have suggested that astrometry can be used to identify and constrain lunar companions within a system (Cabrera and Schneider, 2007; Agol et al., 2015; van Woerkom and Kleisioti, 2024). This is a particularly exciting method given the recent advancements in measuring precise astrometry of exoplanets using long-baseline interferometry. Recently, Lacour et al. (2021) showcased the high-precision astrometric capabilities of the GRAVITY instrument (GRAVITY Collaboration et al., 2017) on the Very Large Telescope Interferometer (VLTI), which can make astrometric measurements with precision of tens of microarcseconds. Building on these capabilities, Winterhalder et al. (2025) demonstrated through simulations and VLTI/GRAVITY observations that moons as small as 0.14 could already be inferred around the nearby (d26.9 pc) AF Lep b system, with future facilities expected to reach Earth-like moons in habitable zones. Complementary work by Wagner et al. (2025) presented a simulation framework for detecting exomoons around imaged giant planets, showing that Earth-mass moons in the habitable zone of Centauri A could be detected with future space- or ground-based facilities. Most recently, Kral et al. (2025) placed constraints on moons orbiting the brown dwarf HD 206893 B with GRAVITY astrometry and detected a tentative candidate. Together, these studies highlight how astrometry is emerging as a promising pathway toward the first confirmed detections of exomoons.
@@
One of the most promising targets for astrometric studies is the Beta () Pictoris system, a young and nearby planetary system that has been extensively studied since the discovery of its circumstellar debris disk in the 1980s (Smith and Terrile, 1984). This system hosts two directly imaged giant planets, Pic b and c, making it ideal for testing astrometric detection techniques. Pic b, located at a semi-major axis of AU, was discovered through direct imaging, while Pic c, at AU, was discovered through RV measurements . Lagrange et al. (2020) combined astrometry and RV data to constrain the orbits and masses of both planets, finding that Pic b has a mass of and Pic c has a well-constrained mass of .
@@
In this work, we use archival data from Lagrange et al. (2020) and Lacour et al. (2020b) to place constraints on the presence of exomoons and binary planets in the Pictoris system using high-precision astrometric measurements from the GRAVITY instrument. While we find no evidence supporting the existence of an exomoon, our analysis establishes strong constraints on the mass and orbital period of any potential companions, providing a framework for applying astrometric techniques to future exomoon searches. Our paper is organized as follows: Section II outlines the instrumentation and data. Section III describes our analysis methods, including the development and application of our model. Section IV presents our results, focusing on the derived mass and orbital constraints. Section V then discusses the implications of our findings, comparing astrometry with complementary techniques and addressing the potential for future studies. We conclude in Section VI.
II Observations
@@
In this work, we used existing archival direct imaging and interferometric observations of Pictoris b and c. Pictoris has been observed for over a decade by numerous instruments, and the precision of astrometric measurements has increased over time. We focused on data taken with the GRAVITY interferometer (GRAVITY Collaboration et al., 2017) and the SPHERE high-contrast imager (Beuzit et al., 2019), provided in Table 1. Following , we excluded earlier, less-precise data from the NACO (NAOS-CONICA) instrument (Lenzen et al., 2003; Rousset et al., 2003). Together, these datasets provide a high-precision foundation for our orbital analysis.
@@ II.2 GRAVITY
We use GRAVITY K band observations of Pictoris b and c . The observations collected by Lagrange et al. (2020) took place on six nights between September 22, 2018, and February 9, 2020. Three of the observations were obtained in the medium-resolution (R = 500) mode of the GRAVITY spectrometer, and two were obtained in the high-resolution mode (R = 4000). On January 7th, both modes were used. Following Lacour et al. (2021), we exclude two of the GRAVITY observations of Lagrange et al. (2020) (MJDs 58738.267 and 58855.198) from our analysis due to poor coherence time and limited exposures. We also include in our analysis additional observations collected by Lacour et al. (2020b) as part of the ExoGRAVITY program (Lacour et al., 2020a) on the night of January 6, 2021 for Pictoris c and the nights of January 7 and August 27, 2021, for Pictoris b. The astrometric calibration from both Lagrange et al. (2020) and Lacour et al. (2021) underwent processing using version 1.5.0 of the ESO GRAVITY pipeline (Lapeyrere et al., 2014).
III Analysis
@@ III.1 Brightness Ratio
III.2 Two Planet Model and Likelihood Function
To fit the SPHERE and GRAVITY data, we constructed a model capable of accurately predicting the positions of Pic b and c over time. To create this initial two-planet model, we adapted code from the Orbitize! Python package to simultaneously calculate the right ascension (RA) and declination (DEC) offsets of both planets with respect to the star. The model also accounts for the astrometric perturbation on Pic b caused by its inner companion, Pic c, following the .
We constructed a log likelihood function to compare the archival data (see Section II) with this model. Maximizing the log likelihood () determines the best-fit orbital parameters for Pic b and c, and exploring the parameter space constrains the plausible parameter space for any orbiting exomoon. to adjust the GRAVITY uncertainties such that:
@@ III.2 Two Planet Model and Likelihood Function
III.3 Two Planet & Moon Model
To model the astrometric signature of a hypothetical dark exomoon orbiting Pic b, we extend the two-planet model described in Section III.2 by adding a Keplerian companion on a bound orbit around Pic b. We used the same log likelihood functions as described in Equations 1 through 4, but include an additional astrometric perturbation term representing the motion of Pic b about the barycenter of the Pic b-moon system. To do this, we used orbitize to generate Keplerian position offsets (RA, DEC) for an exomoon orbiting Pic b, where its gravitational influence shifts the apparent position of Pic b by:
@@ III.3 Two Planet & Moon Model
| (M⊙) | 1.701 | 1.710 | 1.73 |
| Parallax (mas) | 51.39 | 51.40 | 51.44 |
| Error Scale Factor | 1.31 | 1.38 | – |
@@
| Semi-Major Axis (au) | 10.03 | 10.02 | 9.95 |
| Eccentricity | 0.1094 | 0.1072 | 0.106 |
| Inclination (deg) | 88.991 | 88.995 | 88.99 |
| Argument of Periastron (deg) | 212.5 | 211.1 | 203.2 |
| Longitude of Ascending Node (deg) | 31.773 | 31.771 | 31.80 |
| Epoch of Periastron Passage | 0.774 | 0.769 | 0.730 |
@@
| Semi-Major Axis (au) | 2.688 | 2.66 | 2.61 |
| Eccentricity | 0.235 | 0.234 | 0.32 |
| Inclination (deg) | 88.94 | 88.92 | 88.92 |
| Argument of Periastron (deg) | 61 | 57 | 60.8 |
| Longitude of Ascending Node (deg) | 31.18 | 31.21 | 31.07 |
| Epoch of Periastron Passage | 0.825 | 0.82 | 0.708 |
III.4 Fitting
@@
IV Results
IV.1 Constraints on Exomoons Around Pic b
@@ IV.2 Orbital Parameters for Pic b and c
V Discussion
V.1 Constraints on Exomoons and Binary Planets in Context
@@
The moons that we could detect in the Pic system are not like the moons in our own solar system. The Solar System’s most massive moon is Ganymede, which has a mass 50 times smaller than that of Earth, and 1500 times smaller than the lowest mass moon we can rule out at 3 around Pic b. However, the moons and binary planets that we could detect in the Pic system have similar mass ratios to some of the Solar System moons. At long orbital periods, we could detect moons with masses of about 1%, which is smaller than both the Pluto/Charon and Earth/Moon mass ratios. It is worth noting that both the Earth/Moon and Pluto/Charon systems are believed to have formed via collision and capture of lunar material (Denton et al., 2025; Canup and Asphaug, 2001); any moons we could detect might likely have formed similarly in their systems.
V.2 Are we sensitive to moons or binary planets?
@@
Given that we are only sensitive 222. to Neptune-mass companions and larger, it is worth asking: are the objects that we could detect moons, or should they instead be considered binary planets? The classification of a satellite as a moon or a binary planet remains an open question, particularly when considering systems beyond our solar system. A simple way to distinguish moons and binary planets would be to choose a mass ratio between the two orbiting objects above which an orbiting body is a binary planet, and below which it is a moon. Such a strategy has been applied to the IAU definition of a planet (Lecavelier des Etangs and Lissauer, 2022), which states that the mass ratio must be large enough that stable orbits can exist around the L4 and L5 Lagrange points: , or more precisely . If we were to, by extension, apply this definition to distinguish moons from binary planets, our analysis is sensitive to some exomoons around Pic b at orbital periods longer than about 100 days.
@@
However, there is no official IAU definition for a moon, so instead, other criteria have been utilized. According to one commonly cited definition from Stern and Levison (2002), a moon is a satellite with a mass and semi-major axis small enough to keep the center of mass of the moon-planet system within the radius of the primary planet. In contrast, a binary planet is defined as a system in which the companion is sufficiently massive for the center of mass to extend beyond the primary planet. By this definition, we are not sensitive to exomoons around Pic b. We show the Stern and Levison (2002) boundary between moons and binary planets in Figure 4 as a solid green line. This boundary falls slightly below our 1 contour at all orbital periods and scales in a similar manner with orbital period to our sensitivity limits. This similar scaling is expected, since the Stern and Levison (2002) definition of a moon and the amplitude of astrometric signals both depend on the center of mass motion of the planet.
@@
However, both of these definitions have drawbacks. For instance, although Charon formally meets the requirement for a binary planet from both of these two definitions, it is still commonly referred to as a moon of Pluto. Moreover, under the Stern and Levison (2002) definition, the same object at different orbital separations may be either a moon or a binary planet. Given orbital evolution, this implies that an orbiting companion may be both a moon and a binary planet at different times in the system’s history. For instance, while the Earth/Moon system currently meets the Stern and Levison (2002) criterion for a moon, after the Moon’s semimajor axis has increased by about 40% due to tidal evolution, it would by this definition formally become a binary planet. It is also plausible that because planet radii contract over time, a young planet/moon system may evolve into a binary system as the planet radius becomes small enough that it does not encompass the system’s center of mass. Ultimately, a precise distinction between moons and binary planets will likely have to wait until astronomers have a better understanding of the broader population beyond the solar system.
| Orbital Period (d) | 68% (M) | 95% (M) | 99.7% (M) |
| 2.424749 | 1.306464 | 3.461267 | 5.558522 |
| 7.2742474 | 0.853945 | 2.157875 | 3.674100 |
| 12.123745 | 0.485129 | 1.122373 | 2.067631 |
| 16.973244 | 0.336653 | 0.734944 | 1.370374 |
| 21.822742 | 0.274056 | 0.590851 | 1.096990 |
V.3 Prospects for Improved Exomoon Searches in the Future